# The synthetic instrument: From sparse association to sparse causation

### v2

@@ 6 Real data application
The estimator proposed by Onatski (2010) suggests that there are three unobserved latent factors, so we applied our method with $\widehat{q}=3$, using a linear outcome model as the working model. Five genes were found to affect mouse body weight. Under the plurality rule A4, we conclude that the causal effects are identifiable as $\widehat{q}+\widehat{s}=8\ll p=2819$. Specifically, our analysis suggests that increasing the gene expression levels of $Igfbp2$, $Rab27a$, $Dct$, $Ankhd1$, and $Gck$ by one standard deviation leads to changes of $-1.98$, $1.88$, $1.43$, $-1.33$, and $1.17$ grams in mouse body weight, respectively.
We compared our approach with six methods: (i) the Lasso method; (ii) the two-stage regularization (2SR) method (Lin et al., 2015), which leverages SNPs as high-dimensional instrumental variables to estimate the causal effect; (iii) the auxiliary variable method (Miao et al., 2023), which focuses on the 17 genes identified by Lin et al. (2015) as having non-zero effects and uses the five SNPs selected by Lin et al. (2015) as auxiliary variables; (iv) the null variable method (Miao et al., 2023), which assumes that more than half of the 17 genes have zero effects on mouse body weight; (v) the Trim method (Ćevid et al., 2020); and (vi) the IV-Lasso method, as detailed in Section 5. Detailed results for these comparison methods are included in Supplementary Material Section LABEL:sec:comparison. The number of active genes found by these methods, defined as genes that have non-zero effects on mouse body weight, is 87, 17, 4, 2, 4, and 14, respectively. Consistent with the simulation results, the Lasso method identifies many more active exposures than our method, and its selected set includes all genes identified by our approach. Compared to the other methods, both the 2SR and auxiliary variable methods rely on additional SNP information.
@@ 7 Discussion
In this paper, we study how to identify and estimate causal effects with a multi-dimensional treatment in the presence of unmeasured confounding. Our key assumption is a sparse causation assumption, which in many contexts serves as an appealing alternative to the widely adopted sparse association assumption. We develop a synthetic instrument approach to identify and estimate causal effects without the need for collecting additional exogenous information, such as instrumental variables. Our estimation procedure can be formulated as an $\ell_{0}$-optimization problem and can therefore be solved efficiently using off-the-shelf packages.
@@
We have focused on a linear treatment model (1). In a general nonlinear treatment model, where the relationship between treatment $X$ and the unmeasured factor $U$ is nonlinear, nonlinear factor analysis could be used to fit $X=m(U;\alpha)+\epsilon$. However, identifying a function $h(\cdot)$ such that $h(X)$ is independent of $U$ remains a significant challenge. Extending this framework to accommodate nonlinear treatment models is left for future research.
@@ Acknowledgements
We thank Xin Bing, Zhichao Jiang, Wang Miao, Thomas Richardson, James Robins, Dominik Rothenhäusler, and Xiaochuan Shi for their helpful discussions and constructive comments. We also extend our gratitude to the Editor, Associate Editor, and anonymous referees for their valuable and thoughtful input, which has significantly enhanced the quality of this manuscript.

### v3

@@ 6 Real data application
The estimator proposed by Onatski (2010) suggests that there are three unobserved latent factors, so we applied our method with $\widehat{q}=3$, using a linear outcome model as the working model. Five genes were found to affect mouse body weight. Under the plurality rule A4, we conclude that the causal effects are identifiable as $\widehat{q}+\widehat{s}=8\ll p=2819$. Specifically, our analysis suggests that increasing the gene expression levels of $Igfbp2$, $Rab27a$, DctDct, $Ankhd1$, and $Gck$ by one standard deviation leads to changes of 1.98([2.69,1.27])-1.98([-2.69,-1.27]), $1.88([1.26,2.51])$, $1.43([0.86,2.02])$, 1.33([1.90,0.77])-1.33([-1.90,-0.77]), and $1.17([0.69,1.66])$ grams in mouse body weight, respectively. The empirical confidence intervals were constructed using the ivreg function; see Section LABEL:sec:CI_sup for more details.
We compared our approach with six methods: (i) the Lasso method; (ii) the two-stage regularization (2SR) method (Lin et al., 2015), which leverages SNPs as high-dimensional instrumental variables to estimate the causal effect; (iii) the auxiliary variable method (Miao et al., 2023), which focuses on the 17 genes identified by Lin et al. (2015) as having non-zero effects and uses the five SNPs selected by Lin et al. (2015) as auxiliary variables; (iv) the null variable method (Miao et al., 2023), which assumes that more than half of the 17 genes have zero effects on mouse body weight; (v) the Trim method (Ćevid et al., 2020); and (vi) the IV-Lasso method, as detailed in Section 5. Detailed results for these comparison methods are included in Section LABEL:sec:comparison of the supplementary material. The number of active genes found by these methods, defined as genes with non-zero effects on mouse body weight, is 87, 17, 4, 2, 4, and 14, respectively. Consistent with the simulation results, the Lasso method identifies many more active exposures than our method, and its selected set includes all genes identified by our approach. Compared to the other methods, both the 2SR and auxiliary variable methods rely on additional SNP information.
@@ 7 Discussion
In this paper, we study how to identify and estimate causal effects with a multi-dimensional treatment in the presence of unmeasured confounding. Our key assumption is a sparse causation assumption, which in many contexts serves as an appealing alternative to the widely adopted sparse association assumption. We develop a synthetic instrument approach to identify and estimate causal effects without the need to collect additional exogenous information, such as instrumental variables. Our estimation procedure can be formulated as an $\ell_{0}$-optimization problem and can therefore be solved efficiently using off-the-shelf packages.
@@
We have focused on a linear treatment model (1). In a general nonlinear treatment model, where the relationship between treatment $X$ and the unmeasured factor $U$ is nonlinear, nonlinear factor analysis could be used to fit $X=m(U;\Lambda)+\epsilon$. However, identifying a function $h(\cdot)$ such that $h(X)$ is independent of $U$ remains a significant challenge. Extending this framework to accommodate nonlinear treatment models is left for future research.
@@ Acknowledgements
We thank Xin Bing, Zhichao Jiang, Wang Miao, Thomas Richardson, James Robins, Dominik Rothenhäusler, and Xiaochuan Shi for their helpful discussions and constructive comments. We also extend our gratitude to the Editor, the Associate Editor, and the anonymous referees for their valuable and thoughtful input, which has significantly improved the quality of this manuscript.